# T2S: High-resolution Time Series Generation with Text-to-Series Diffusion Models

### v1

@@ 1 Introduction
Time series generation (TSG) enables the creation of high-quality data in scenarios with limited available datasets, thus serving as a way to simulate diverse, multimodal temporal dynamics, which offers significant value in real-world applications. Generating required objects from text (Text-to-X) is a research area that meets human needs and has potential at the present time. Driven by the success of diffusion models, Text-to-X generation has made remarkable strides in domains such as image generation Rombach et al. (2022); Esser et al. (2024), video generation Girdhar et al. (2023), and speech processing Le et al. (2024). Diffusion models refine noisy data progressively through a learned process, ultimately producing high-fidelity outputs. While Text-to-X generation using diffusion models has been extensively explored in vision and audio data, their application to time series generation is still in its early stages.
Existing studies Yang et al. (2024) on TSG with diffusion models can be classified into three types based on the conditioning information. i) Label-based condition. A class-conditioned diffusion method for generating synthetic EEG signals is introduced in Sharma et al. (2023). ii) Temporal-based condition. Sensor data synthesis using statistical adjustments is explored in Zuo et al. (2023). Although class- and temporal-conditioned generation is well discussed Yuan and Qiao (2024); Liu et al. (2024b), they are less flexible than text-based approaches Gao et al. (2024). iii) Text-based condition. Fu et al. (2024) uses domain-specific metadata (e.g., location, weather) to generate synthetic energy data, while Wang (2024) and Lai et al. (2025); Alcaraz and Strodthoff (2023) apply textual conditioning to synthesize sales and clinical ECG data, respectively. However, these methods are often domain-specific, their metadata captions can not support high resolution general alignment between time series and fine-grained captions. Despite the progress made in applying diffusion models to time series modeling, two significant challenges remain in this domain. First, the scarcity of high-resolution general-propose text-time series caption datasets limits progress in Text-to-Time Series (T2S) generation, while existing datasets are domain-specific (e.g., healthcare Johnson et al. (2023), economics Cortis et al. (2017)). Second, existing TSG models Yuan and Qiao (2024); Desai et al. (2021) typically require separate training for datasets of different lengths within each domain, making it challenging to develop length-arbitrary T2S models. Most approaches rely on predefined sequence lengths tied to the training data, limiting their ability to generalize. Since real-world time series exhibit inherent variability in length due to factors such as data collection frequency or system-specific temporal dynamics, the need for length-specific training significantly hinders the scalability and practicality of these models.
@@ 2.1 Problem Definition and Notation
A point-level description $\mathbf{T}_{p}$ provides semantic annotations for individual time points within the time series $\mathbf{x}$. Each $\mathbf{T}_{p}^{(j)}$ corresponds to the $j$-th time point, where $j\in[1,\ldots,L]$, providing a fine-grained guidance for time series $\mathbf{x}$.
@@
A fragment-level description $\mathbf{T}_{f}$ provides semantic annotations for non-overlapping and contiguous fragments of the time series $\mathbf{x}$. Each $\mathbf{T}_{f}^{(j)}$ corresponds to the $j$-th fragment, where the length $|\mathbf{T}_{f}^{(j)}|$ is arbitrary and determined by the specific structure of time series $\mathbf{x}$.
@@
Given a text-time series dataset $D=\{(\mathbf{x}^{(i)},\mathbf{T}_{*}^{(i)})\}_{i=1}^{N}$ of $N$ samples, the task of T2S generation is to learn a generative model $G$ that maximizes the conditional probability $P_{G}(\mathbf{x}\mid\mathbf{T}_{*})$, where $\mathbf{T}_{*}$ is the textual guidance provided at one of the levels ($\mathbf{T}_{p},\mathbf{T}_{f},\mathbf{T}_{i}$). The generated time series $\mathbf{x}$ must semantically align with the textual guidance $\mathbf{T}_{*}$.
@@ 2.2 TSFragment-600K
The proposed T2S generation task leverages captions at three granularity levels: point, fragment, and instance. While datasets for point-level Liu et al. (2024a) and instance-level Kawaguchi et al. (2025) descriptions are readily available, fragment-level descriptions remain an underexplored area. Fragment-level descriptions strike a balance between the granularity of point-level annotations, which may overlook broader temporal patterns, and the holistic nature of instance-level descriptions, which might obscure local dependencies. By encapsulating local temporal trends while preserving meaningful contextual relationships, fragment-level descriptions provide an ideal framework for evaluating the generative capabilities of T2S models.
@@
Specifically, we propose a novel generation pipeline to construct fragment-level captions for time series data. First, a univariate time series $\mathbf{x}$ is segmented into $k$ non-overlapping fragments,with each fragment fragment $\mathbf{x}^{(j)}\in\mathbb{R}^{l_{j}}$ representing a contiguous temporal segment for which textual captions are generated. Second, a seed-based prompting strategy is designed to capture high-quality captions of fragments. To accurately capture temporal dynamics, human experts curate high-quality GPT-4o descriptions for a subset of fragments, which serve as representative seed prompts. These prompts guide GPT-4o-mini in generating concise, consistent, and semantically rich captions for all fragments. A token limit $\langle\boldsymbol{M}\rangle$ is applied to ensure a balance between informativeness and brevity.
Finally, five candidate captions are generated for each time series sample, and their embeddings are computed using text-embedding-3-small. We ensure the quality of generated textual captions by leveraging cosine similarity among their embeddings. Each caption is scored based on the average similarity of its embedding with others, and the one with the highest score is selected as the optimal text-time series pair, ensuring semantic alignment and coherence. Using this pipeline, we generate reliable fragment-level descriptions for eight classical time series datasets across diverse domains, including energy consumption, financial exchange rates, traffic, air quality, and meteorological variables. The resulting dataset, TSFragment-600K, comprises over $600,000$ samples with corresponding captions.
@@ 3 Methodology
  • T2S Diffusion Transformer (T2S-DiT): T2S-DiT facilitates high-resolution alignment between captions and the temporal latent space. It employs flow matching Liu et al. (2022) as the diffusion backbone, the diffusion transformer module Peebles and Xie (2023) as the denoiser, and integrates text conditions into the denoiser via adaptive layer normalization.
@@ 3.1 T2S Diffusion Model
During sampling, pure noise $\mathbf{z}_{0}\sim\mathcal{N}(0,1)$ is iteratively denoised to obtain $\mathbf{\hat{z}}_{1}$ by solving the ordinary differential equation with the well-trained denoiser $\mathbf{u}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{C}\right)$.
T2S adopts a classifier-free guidance framework, which does not rely on explicit class labels for conditioning. Unlike traditional class-based conditional generation methods, which use fixed condition categories to guide the generation process, classifier-free guidance provides a more flexible and effective way to balance unconditional and conditional generation. During training, the method randomly sets the condition to zero by a random ratio. During inference, the model first performs conditional generation, followed by unconditional generation, and then combines the results using a guidance scale $\delta$. The formula is:
@@
where $\mathbf{u}_{\theta}(\mathbf{z}_{t},t,\mathbf{C})$ is the noise estimate with input $\mathbf{z}_{t}$ and condition $\mathbf{C}$, while $\mathbf{u}_{\theta}(\mathbf{z}_{t},t)$ is the noise estimater without condition.
@@ 3.2 Length-adaptive VAE
T2S generation should support arbitrary-length generation to meet real-world application demands. For example, website user activity analysis and medical monitoring. However, previous works Lai et al. (2025) typically assume a fixed length, limiting their applicability. To address this limitation, we propose a pretrained LA-VAE, enabling the modeling and generation of time series with arbitrary lengths. In this study, we mixed data with lengths of 24, 48, and 96, training them in a unified framework. During sampling, arbitrary-length data can be generated within a specified range. Given an input time series $\mathbf{x}$, the LA-VAE encoder transforms $\mathbf{x}$ into a latent representation $\mathbf{h}$. This latent representation $\mathbf{h}$ is subsequently upsampled to a fixed-size latent embedding $\mathbf{z}=\operatorname{Up-Samp}(\mathbf{h}_{t})$, which serves as the input to the diffusion model. Through the diffusion process, a refined latent embedding $\hat{\mathbf{z}}_{t}$ is generated. A downsampling operation is then applied to $\hat{\mathbf{z}}_{t}$, yielding the latent vector $\hat{\mathbf{h}}=\operatorname{Down-Samp}(\hat{\mathbf{z}}_{t})$. Finally, the VAE decoder reconstructs $\hat{\mathbf{h}}$ into a time series with the original length. This process enables the handling of variable-length time series within a unified framework.
@@ 4.1 Experimental Settings
  • $\bullet$
    Point-Level Dataset: The Time-MMD dataset Liu et al. (2024a) links individual time series points with corresponding textual news, consisting of 23,618 data points across six domains, including Climate, Economy, Energy, Health, and Social Goods. We adapted the dataset by concatenating each time series point with its associated text.
  • $\bullet$
    Fragment-Level Dataset: TSFragment-600K pairs time series data with captions across seven domain, including Electricity, ETT (four subsets), Exchange, and Traffic Wu et al. (2021). It consists of over 600,000 samples.
  • $\bullet$
@@
    $\operatorname{WAPE}(y,\hat{y})=\frac{\sum_{i\in\Omega}\left|y_{i}-\hat{y}_{i}\right|}{\sum_{i\in\Omega}\left|y_{i}\right|}$ (11)
@@
    $\operatorname{MRR@10}=\frac{1}{|\Omega|}\sum_{i\in\Omega}\frac{1}{\text{rank}_{i}}$ (12)
@@ 4.3 Performance Comparison on Point and Instance-Level Descriptions (RQ2)
Table 2 presents the performance comparison at the point and instance levels across seven datasets. These evaluations assess models’ abilities to capture fine-grained temporal annotations and generate coherent global patterns. T2S outperforms all baselines across three metrics. It consistently achieves the lowest MSE values in 17 out of 18 entries, surpassing DiffusionTS and TimeVAE, which struggle to capture fine-grained temporal variations. Similarly, T2S secures the best WAPE scores in 16 out of 18 entries, demonstrating its robustness. It is worth noting that on the instance-level SUSHI dataset, T2S achieves the highest MRR@10 value of 0.314. In terms of WAPE and MSE, T2S outperforms the zero-shot models but performs slightly worse than DiffusionTS. In summary, T2S demonstrates state-of-the-art performance by effectively balancing fine-grained precision with high-level semantic understanding.

### v2

@@ 1 Introduction
Time series generation (TSG) enables the creation of high-quality data in scenarios with limited available datasets, thus serving as a way to simulate diverse, multimodal temporal dynamics, which offers significant value in real-world applications. Generating required objects from text (Text-to-X) is a research area that meets human needs and has potential at the present time. Driven by the success of diffusion models, Text-to-X generation has made remarkable strides in domains such as image generation Rombach et al. (2022); Esser et al. (2024), video generation Girdhar et al. (2023), and speech processing Le et al. (2024). Specifically, diffusion models refine noisy data progressively through a learned process, ultimately producing high-fidelity outputs. While Text-to-X generation using diffusion models has been extensively explored in vision and audio data, their application to time series generation is still in its early stages.
Existing studies Yang et al. (2024) on TSG with diffusion models can be classified into three types based on the conditioning information. i) Label-based condition. A class-conditioned diffusion method for generating synthetic EEG signals is introduced in Sharma et al. (2023). ii) Temporal-based condition. Sensor data synthesis using statistical adjustments is explored in Zuo et al. (2023). Although class- and temporal-conditioned generation is well discussed Yuan and Qiao (2024); Liu et al. (2024b), they are less flexible than text-based approaches Gao et al. (2024). iii) Text-based condition. Fu et al. (2024) uses domain-specific metadata (e.g., location, weather) to generate synthetic energy data, while Wang (2024) and Lai et al. (2025); Alcaraz and Strodthoff (2023) apply textual conditioning to synthesize sales and clinical ECG data, respectively. However, these methods are often domain-specific, their metadata captions can not support high resolution general alignment between time series and fine-grained captions.
Despite the progress made in applying diffusion models to time series modeling, two significant challenges remain in this domain. First, the scarcity of high-resolution general-propose text-time series caption datasets limits progress in Text-to-Time Series (T2S) generation, while existing datasets are domain-specific (e.g., healthcare Johnson et al. (2023), economics Cortis et al. (2017)). Second, existing TSG models Yuan and Qiao (2024); Desai et al. (2021) typically require separate training for datasets of different lengths within each domain, making it challenging to develop length-arbitrary T2S models. Most approaches rely on predefined sequence lengths tied to the training data, limiting their ability to generalize. Since real-world time series exhibit inherent variability in length due to factors such as data collection frequency or system-specific temporal dynamics, the need for length-specific training significantly hinders the scalability and practicality of these models.
@@ 2.1 Problem Definition and Notation
A point-level description $\mathbf{T}_{p}$ provides semantic annotations for individual time points within the time series $\mathbf{x}$. Each point-level description $\mathbf{T}_{p}^{(j)}$ corresponds to the $j$-th time point, where $j\in[1,\ldots,L]$, providing a fine-grained guidance for each point in time series $\mathbf{x}$.
@@
A fragment-level description $\mathbf{T}_{f}$ provides semantic annotations for non-overlapping and contiguous fragments of the time series $\mathbf{x}$. Each fragment-level description $\mathbf{T}_{f}^{(j)}$ corresponds to the $j$-th fragment, where the length $|\mathbf{T}_{f}^{(j)}|$ is arbitrary and determined by the specific structure of time series $\mathbf{x}$.
@@
Given a text–time series dataset $D=\{(\mathbf{x}^{(i)},\mathbf{T}_{*}^{(i)})\}_{i=1}^{N}$ with $N$ samples, the task of T2S generation aims to learn a generative model $G$ that maximizes the conditional probability $P_{G}(\mathbf{x}\mid\mathbf{T}_{*})$, where $\mathbf{T}_{*}$ is the textual guidance provided at one of the following levels ($\mathbf{T}_{p},\mathbf{T}_{f},\mathbf{T}_{i}$). The generated time series $\mathbf{x}$ must semantically align with the textual guidance $\mathbf{T}_{*}$.
@@ 2.2 TSFragment-600K
The proposed T2S generation task leverages textual captions at three granularity levels: point, fragment, and instance. While datasets for point-level Liu et al. (2024a) and instance-level Kawaguchi et al. (2025) descriptions are readily available, fragment-level descriptions remain an underexplored area. Fragment-level descriptions strike a balance between the granularity of point-level annotations, which may overlook broader temporal patterns, and the holistic nature of instance-level descriptions, which might obscure local dependencies. By encapsulating local temporal trends while preserving meaningful contextual relationships, fragment-level descriptions provide an ideal framework for evaluating the generative capabilities of T2S models.
@@
Specifically, we propose a novel generation pipeline to construct fragment-level captions for time series data. First, a univariate time series $\mathbf{x}$ is segmented into $k$ non-overlapping fragments,with each fragment $\mathbf{x}^{(j)}\in\mathbb{R}^{l_{j}}$ as a contiguous temporal segment for which textual captions are generated. Second, a seed-based prompting strategy is designed to capture high-quality captions of fragments. To accurately capture temporal dynamics, human experts curate high-quality GPT-4o descriptions for a subset of fragments, which serve as representative seed prompts. These prompts guide GPT-4o-mini in generating concise, consistent, and semantically rich captions for all fragments. A token limit $\langle\boldsymbol{M}\rangle$ is applied to ensure a balance between informativeness and brevity.
Finally, five candidate captions are generated for each time series sample, and their embeddings are computed using text-embedding-3-small. We ensure the quality of generated textual captions by leveraging cosine similarity among their embeddings. Each caption is scored based on the average similarity of its embedding with others, and the one with the highest score is selected as the optimal text-time series pair, ensuring semantic alignment and coherence. Using this pipeline, we generate reliable fragment-level descriptions for eight classical time series datasets across diverse domains, including energy consumption, financial, exchange rates, traffic, air quality, and meteorological variables. The resulting fragment-level dataset, TSFragment-600K, comprises over $600,000$ samples with corresponding captions.
@@ 3 Methodology
  • T2S Diffusion Transformer (T2S-DiT): T2S-DiT facilitates high-resolution alignment between captions and the temporal latent space. It employs flow matching Liu et al. (2022) as the diffusion backbone, the diffusion transformer module Peebles and Xie (2023) as the denoiser. Within this denoiser, textual information is integrated with the input features through adaptive layer normalization
@@ 3.1 T2S Diffusion Model
During sampling, pure noise $\mathbf{z}_{0}\sim\mathcal{N}(0,1)$ is iteratively denoised to obtain $\mathbf{\hat{z}}_{1}$ by solving the ordinary differential equation (ODE) with the well-trained denoiser $\mathbf{u}_{\theta}\left(\mathbf{z}_{t},t,\mathbf{C}\right)$. T2S adopts a classifier-free guidance framework, which does not rely on explicit class labels for conditioning. Unlike traditional class-based conditional generation methods, which use fixed condition categories to guide the generation process, classifier-free guidance provides a more flexible and effective way to balance unconditional and conditional generation. During training, the method randomly sets the condition to zero by a random ratio. During inference, the model first performs conditional generation, followed by unconditional generation, and then combines the results using a guidance scale $\delta$. The formula is:
@@
where $\mathbf{u}_{\theta}(\mathbf{z}_{t},t,\mathbf{C})$ is the noise estimater with input $\mathbf{z}_{t}$ and condition $\mathbf{C}$ and $\mathbf{u}_{\theta}(\mathbf{z}_{t},t)$ is the conditioned noise estimater .
@@ 3.2 Length-adaptive VAE
T2S generation should support arbitrary-length generation to meet real-world application demands. For example, website user activity analysis and medical monitoring. However, previous works Lai et al. (2025) typically assume a fixed length, limiting their applicability. To address this limitation, we propose a pretrained LA-VAE, enabling the modeling and generation of time series with arbitrary lengths. In this study, we mixed data with lengths of 24, 48, and 96, then trained them in a unified framework. During sampling, arbitrary-length data can be generated within a specified range. Given an input time series $\mathbf{x}$, the LA-VAE encoder transforms $\mathbf{x}$ into a latent representation $\mathbf{h}$. This latent representation $\mathbf{h}$ is subsequently upsampled to a fixed-size latent embedding $\mathbf{z}=\operatorname{Up-Samp}(\mathbf{h}_{t})$, which serves as the input to the diffusion model. Through the diffusion process, a refined latent embedding $\hat{\mathbf{z}}_{t}$ is generated. A downsampling operation is then applied to $\hat{\mathbf{z}}_{t}$, yielding the latent vector $\hat{\mathbf{h}}=\operatorname{Down-Samp}(\hat{\mathbf{z}}_{t})$. Finally, the VAE decoder reconstructs $\hat{\mathbf{h}}$ into a time series with the original length. This process enables the handling of variable-length time series within a unified framework.
@@ 4.1 Experimental Settings
  • Point-Level Dataset: The Time-MMD dataset Liu et al. (2024a) links individual time series points with corresponding textual news, consisting of 23,618 data points across six domains, including Climate, Economy, and Social Goods. We adapted the dataset by concatenating each time series point with its associated text.
  • Fragment-Level Dataset: TSFragment-600K pairs time series data with captions across seven domain, including Electricity, ETT, Exchange, and Traffic Wu et al. (2021). It consists of over 600,000 samples.
@@
    $\operatorname{WAPE}(y,\hat{y})=\frac{\sum_{i\in\Omega}\left|y_{i}-\hat{y}_{i}\right|}{\sum_{i\in\Omega}\left|y_{i}\right|},$ (11)
@@
    $\operatorname{MRR@10}=\frac{1}{|\Omega|}\sum_{i\in\Omega}\frac{1}{\text{rank}_{i}},$ (12)
@@ 4.3 Performance Comparison on Point and Instance-Level Descriptions (RQ2)
Table 2 presents the performance comparison at the point and instance levels across seven datasets. These evaluations assess models’ abilities to capture fine-grained temporal annotations and generate coherent global patterns. T2S outperforms all baselines across three metrics. It consistently achieves the lowest MSE values in 17 out of 18 entries, surpassing DiffusionTS and TimeVAE, which struggle to capture fine-grained temporal variations. Similarly, T2S secures the best WAPE scores in 16 out of 18 entries, demonstrating its robustness. T2S achieves the top MRR@10 score of 0.314 on the instance-level SUSHI dataset, its WAPE and MSE results show an advantage over zero-shot models, only a slight underperformance compared to DiffusionTS in certain scenarios. In summary, T2S demonstrates state-of-the-art performance by effectively balancing fine-grained precision with high-level semantic understanding.